# Low-metallicity massive single stars with rotation

### v2

@@ Abstract
Context. Chemically-homogeneously evolving stars have been proposed to account for several exotic phenomena, including gravitational-wave emissions, gamma-ray bursts and certain types of supernovae.
@@
Results. Adding two new models computed with PoWR, we report (i) these stars to follow a unique sequence of classes: O WN WO (i.e. without ever being WC). From our population synthesis with standard assumptions, we predict that (ii) the source of the UV C IV 1550 Å and other emission bumps is a couple dozen WO-type Wolf–Rayet stars (not WC as previously assumed) which are the result of chemically-homogeneous evolution, while these, combined with the rest of the O-star population, account for the high He II ionizing flux and the spectral hardness. Contrasting our results against published optical and UV data from the literature and accounting for different aperture sizes and spatial regions probed by the observations, we find that (iii) our models are highly consistent with existing measurements.
Conclusions. Since our “massive Pop II stars” might just as well exist in early star-forming regions, our findings have implications for upcoming JWST surveys: first galaxies in the high-redshift Universe may also experience the extra contribution of UV photons and the kind of exotic explosions that chemically-homogeneous stellar evolution predicts. Given that our results apply for binary populations too as long as the same fraction (10%) of the systems evolves chemically homogeneously, we conclude that the stellar progenitors of gravitational waves may very well exist today in I Zw 18.
@@
Galaxies: dwarf—Stars: massive—Stars: Wolf-Rayet—(Cosmology:) dark ages, reionization, first stars—Gravitational waves1 Introduction
@@
Dwarf starburst galaxies form stars at a high rate, and thus harbour large populations of metal-poor massive stars (e.g. Zhao et al. 2013). Despite the recent burgeoning of interest in these stars’ life (Weisz et al. 2014; Izotov et al. 2016; Schootemeijer & Langer 2018; Evans et al. 2019; Garcia et al. 2019; Groh et al. 2019; Senchyna et al. 2019; Franeck et al. 2022; Szécsi et al. 2022; Gull et al. 2022; Lahén et al. 2023) and death as exotic explosions (Perley et al. 2016; Aguilera-Dena et al. 2018; Stevenson et al. 2019; Agrawal et al. 2022), and ‘afterlife’ as—potentially merging—compact objects (Eldridge & Stanway 2016; Marchant et al. 2016; de Mink & Mandel 2016; Mandel & de Mink 2016; Marchant et al. 2017; Vigna-Gómez et al. 2018; Eldridge et al. 2019; Romagnolo et al. 2022), theoretical models are far from being properly tested observationally (Mintz et al. 2025).
@@
One convenient testbed for constraining our models is local dwarf star-forming galaxies. These objects typically display low metallicities and high star-formation rates (Searle & Sargent 1972; Hunter & Thronson 1995; Vílchez & Iglesias-Páramo 1998; Izotov & Thuan 2002, 2004; Thuan & Izotov 2005; Vaduvescu et al. 2007; Shirazi & Brinchmann 2012; Annibali et al. 2013; Kehrig et al. 2013, 2018, 2021; Arroyo-Polonio et al. 2024; Hirschauer et al. 2024), meaning that they are local equivalents for high-redshift star-forming regions. Studying them can shed light to the nature of metal-poor massive-star populations which, presumably, contribute to the structure of the galaxy with their radiative, chemical and mechanical feedback.
@@
In the blue compact dwarf galaxy I Zw 18 in particular, a high amount of photoionization has been derived from nebular line observations: Kehrig et al. (2015b) reported a He II ionizing flux of Q 1.331050 photons s-1 (see their Sect. 3.2). I Zw 18 has two large clusters, a North-Western cluster and a South-Wastern cluster. Out of these two, only the North-Western cluster was spacially associated with this nebular emission; the total stellar mass of this region is about 300 000 M⊙ (Izotov & Thuan 1998; Lebouteiller et al. 2013). The high He II flux was puzzling because it could not be accounted for by the ionization from conventional ionizing sources including standard hot massive stars (Kehrig et al. 2015b).
@@
The hottest and most luminous massive stars are Wolf–Rayet (WR) stars: they are even stronger ionizing sources than O stars (Hamann et al. 2015). Due to their optically thick stellar winds, WR stars also emit broad emission lines. One such line is the UV-line C IV 1550 Å which has indeed been identified in the spectrum of I Zw 18 (Lebouteiller et al. 2013): in the North-Western cluster, a line luminosity between L 2.25.51037 erg s-1 has been measured by Brown et al. (2002). The problem is that according to classical stellar theory, the number of WR stars derived from the carbon-line luminosity cannot account for the high He II emission observed. As Kehrig et al. (2015b) reports, to explain L (which they take to be to be 4.671037 erg s-1, cf. their Sect. 4.1), about 9 classical Wolf–Rayet stars of type WC (with standard WC-properties given by Crowther & Hadfield 2006, also cf. López-Sánchez & Esteban 2010a, b and Sander 2022) are required in the region: and the ionizing flux of only 9 of them is 50 times lower than Q.
@@
Therefore, other candidates were suggested as the source of ionization. One option is that a population of peculiar (nearly) metal-free massive stars (Pop III-star siblings/Pop III-like stars Kehrig et al. 2015b, a; Heap et al. 2015) formed recently from some leftover primordial gas in this galaxy. Pop III stars are expected to be hotter than metal-poor stars without developing carbon emission lines (Yoon et al. 2012). This hypothesis relies on I Zw 18 keeping, or somehow obtaining, pockets of primordial gas that have just recently started to form massive stars.
@@
Another scenario was presented by Péquignot (2008), Lebouteiller et al. (2017) and Schaerer et al. (2019), in which a population of X-ray binaries produce the ionizing radiation. However, Kehrig et al. (2021) found that neither a high-mass X-ray binary, nor the soft X-ray photons observed in the galaxy can account for the bulk of the nebular HeII emission; and Senchyna et al. (2020) support this conclusion. In turn, Oskinova & Schaerer (2022) suggest cluster winds and superbubbles as an alternative explanation.
@@
Here we study yet another possibility. We simulated stellar evolutionary models of single massive stars with the exact composition of I Zw 18 in Szécsi et al. (2015, hereafter Paper I), using state-of-the-art physics. We found that these “massive Pop-II stars” mostly evolve the classical way (similar to how metal-rich massive stars in our Galaxy do) as long as their rotational rate is moderate. But those models that rotate fast follow another path: they evolve chemically homogeneously, and become hot and luminous stars—similar to Wolf-Rayet stars but without the high wind-mass-loss and thus without the broad emission features during most of their lives. To analyse these objects further, we computed synthetic spectra for them in Kubátová et al. (2019, hereafter Paper II). We dubbed them Transparent Wind UV-Intense stars (in short TWUIN stars) to distinguish them from traditional Wolf–Rayet stars with optically thick winds.
@@
Below we investigate whether a metal-poor massive star population based on our evolutionary models from Paper I and our spectral models from Paper II can simultaneously account for both the observed He II ionizing flux and the presence/absence of emission lines in I Zw 18. Section 2 explains our modelling approach: from evolutionary models (Sect. 2.1) to synthetic spectra (Sect. 2.2, including the addition of new PoWR spectra and the newly established classification sequence for chemically-homogeneously evolving stars) to population synthesis (Sect. 2.3) and emission line synthesis (Sect. 2.4). Section 3 compares our theoretical predictions to observations in terms of ionizing radiation and C IV 1550 Å emission. Then in Section 4, we dive into the existing literature to figure out if other emission lines observed so far in I Zw 18 (both optical and UV) support or contradict our theory. Section 5 discusses our explanation’s place amongst alternative ones as well as the necessary caveats. In Section 6 we conclude that, as far as the currently available observational data allows, I Zw 18 may—similarly to high-redshift galaxies (Liu et al. 2025)—very well harbor chemically-homogeneously evolving stars.
@@ 2.1 The models: predicting O and WO stars
Metal-poor massive single stars have been modelled using the ‘Bonn’ stellar evolution code in Paper I with the measured composition of I Zw 18 (totalling in a metallicity of Z=0.0002111 Metallicities of dwarf galaxies are usually inferred from their nebular O III lines, which can be problematic since the ratio of iron to oxygen (and all other elements) does not necessarily follow solar patterns—on this, see Sect. 2.2.1 of Chruslinska & Nelemans (2019). Different works report different metallicies for I Zw 18, e.g. 1/40 Z⊙ in Kehrig et al. (2016) and 1/30 Z⊙ in Izotov & Thuan (1998); this may depend on e.g. the calibration used to translate O III equivalent width into abundances. Paper I modelled I Zw 18 relying on element abundances reported in Lebouteiller et al. (2013) corresponding to 1/50 Z⊙. , or [Fe/H] 1.7) and standard input physics. Mass-loss rate prescription from Hamann et al. (1995) scaled down by a factor of 10 and Z-scaling from Vink et al. (2001) was used for the hot helium-star phase, which yields consistent predictions with Nugis & Lamers (2000, for details, see and ). These evolutionary sequences predicted that fast-rotating stars evolve directly towards hot surface temperatures (so-called chemically-homogeneous evolution) and thus serve as potential sources of photoionization in metal-poor dwarf galaxies. Applying simple estimations for the wind optical depth (following Langer 1989) and black body radiation, Paper I showed that during the long-lived main-sequence (core-hydrogen-burning) phase, these hot stars are expected to have transparent winds and therefore should not develop emission lines in their spectra (hence the name TWUIN stars). Paper I estimated that a population of them with a regular mass function can explain the observationally derived He II ionizing flux, Q, in I Zw 18. However, these results were rather crude; a more reliable conclusion could be drawn from synthetic spectra created especially for these stars.
@@
In Paper II, we computed such spectra, again with the composition of I Zw 18, using the Potsdam Wolf-Rayet (PoWR) stellar atmosphere code (Gräfener et al. 2002; Hamann & Gräfener 2003, 2004; Šurlan et al. 2013; Sander et al. 2015; Sander & Vink 2020). We found that chemically-homogeneously evolving stars in the main-sequence phase indeed do not develop emission lines, just as suggested in Paper I. According to standard stellar classification presented in Figure 1, these TWUIN stars, if observed, would be classified as early O type stars (cf. also Table 4 & Appendix A of Paper II). As for the post-main sequence, the evolutionary behavior during this phase was not included into Paper I but was later published by Szécsi et al. 2022 (their grid called ‘IZw18-CHE’) and also studied in Paper II. We found that during core-helium burning, chemically-homogeneously evolving stars do actually develop strong emission lines: they were classified as WR stars of class WO. Fig. 1 summarizes these findings in a HR diagram, and Sect. 2.2 provides a deeper look into the post-main sequence phase.
@@
The spectra published in Paper II were created for three masses (Mini 20, 59 and 131 M⊙, all with high initial rotational rate of 300-600 km-1), five evolutionary stages (four on the main-sequence and one post-main-sequence, see also Sect. 2.2), two mass-loss rates (a nominal one taken from the evolutionary models, and a reduced one; actual values listed in Table 2 of Paper II) and two types of assumptions about wind clumping (a smooth wind and a clumped wind with clumping factor D 10). The mass-loss rate values are supported by those that Hainich et al. (2015) empirically predicted for the same stars. Additionally, more realistic mass-loss rates of these stars were calculated by Abdellaoui (2023) using the VH-1 code which allows multidimensional treatment of the wind. The obtained results (listed in their Table 6.3) are somewhere between our nominal and reduced values.
@@
Fig. 2 shows the spectral energy distribution (SED) of some of our models (those with nominal mass-loss rates and smooth wind), as well as the corresponding black body spectrum for context (without applying wind optical depth correction from Langer 1989). As Fig. 2 attests, the SED of these stars do not follow the distribution of black body radiation, even though the total amount of ionizing photons in the three main ionizing continua (Lyman, He I and He II continua) are on the same order of magnitude as in the (uncorrected) black body.
@@
Having these models allows us to build a synthetic population, including spectral synthesis of various emission lines. Before we do that (Sect. 2.3), we present two new model spectra to properly account for the post-main sequence, a phase crucially relevant for our goals.
@@ 2.2 New PoWR spectra: WN and WO stars (no WC)
In our Paper II, the post-main sequence phase was accounted for by models with central helium mass fraction of YC 0.1, that is, nearly the end of core-helium burning. The spectra were classified as type WO according to the standard spectral classification scheme (for details on this, see Table 4 & Appendix A of Paper II, as well as our Fig. 1), and showed very prominent emission lines in both the optical and the UV regimes. However, according to Figure 3, the evolutionary model predicts significant change in the surface abundances while core-helium burning is in progress, with the YC 0.1 stage being quite different from earlier times. Therefore, before using these late-stage models in our population synthesis (Sect. 2.3), we should justify at which extent they are representative of the entirety of the post-main sequence, and discuss their evolutionary behaviour and spectral classification.
@@
To this end, we computed two more spectral models using PoWR for YC 0.8 and YC 0.5 of a very massive star with initial mass Mini 131 M⊙ (actual mass around 93106 M⊙, see the caption). The computational set-up was the same as in Paper II. We apply nominal mass-loss rates taken from the evolutionary model (Szécsi et al. 2015; Szécsi 2016; Szécsi et al. 2022), and smooth winds (clumping factor D 1); we discuss in Sect. 5.2 that these assumptions are actually quite robust. Figure 4 compares the optical band for these spectra, while Figure 3 summarizes what we learned in terms of stellar classification.
@@
Applying the rudimentary scheme of classification defined by Georgy et al. (2012) for evolutionary models, Szécsi (2016, in their Sect. 4.5.1) suggested these stars to be of class WN, WC and close-to-WO, respectively (Fig. 3). However, the PoWR spectra tell a different story. The standard classification scheme (working with spectral line intensity222This classification scheme follows the footsteps of the Harvard-system. The Harvard system actually included WR-stars, calling them Oa (now WC), Ob (now WNE), Oc (now WNL), cf. Cannon & Pickering (1916), but efforts like Payne (1930), Edlen (1932), and Beals (1933) helped shape the version adopted by the IAU in 1938, which was modernized in the 1960s by Hiltner, Schild and Lindsey Smith, with WO stars first introduced by Barlow & Hummer (1982). Today, we rely on emission line intensities to assign classes to WR-stars, as summarized in Appendix A of Paper II, although switching between subtypes is not as clear as for the absorption-line based classes OBAFGKM and the usage of specific primary and secondary criteria is also less sharp. as summarized by Appendix A of Paper II) yields the following classes: WN2 for YC 0.8, WO 2 or WO 1 (consistent with both) for YC 0.5, and WO 1 for YC 0.1. Based on what we learned in Paper II (especially their Figs. 4 and B6), we expect that reducing the mass loss rate and/or increasing the wind clumping may shift these spectra towards different subtypes, that is, towards different line intensities—but not towards a different class. In other words, varying mass loss and clumping won’t make any of these models a WC star.
@@
We conclude that chemically-homogeneously evolving stars at I Zw 18 metallicity would be seen, after spending their main-sequence as O stars, as early-type WN stars at the beginning, and WO stars in the majority of their post-main sequence lifetimes. This is surprising since typical evolution calculations for single stars predict the WO stage to be a rather late, short-term phase, preceded by a more lengthy WC phase (see, e.g., Groh et al. 2014). The typical sequence of classes read:
O RSG/LBV WN WC WO
@@
Chemically-homogeneous evolution does not follow that picture, as these stars complete their main-sequence as very hot O stars (Sect. 2.1 and Fig. 1) and then proceed to be WN and WO after that, a completely new sequence of classes:
O WN WO
@@
According to Fig.3, if a WC-phase arises at all between WN and WO, it won’t last longer than about 1-2% of the total stellar lifetime.
@@
Finding our stars to be WO instead of WC is in accordance with Tramper et al. (2013) who concluded that WO stars may be the high-temperature and high-luminosity extensions of WC stars. The WO spectral type does not necessarily mean more oxygen than carbon: in our mid-pMS model, the carbon abundance is actually higher than the oxygen one (as seen in Fig. 3). The same was found for the handful of LMC WO stars analysed in Aadland et al. (2022). So it is not about abundance but about temperature: carbon is ionized into C V and thus the otherwise prominent 1550 and 5808 features are weak or absent.
@@
While WC and WO may not necessarily represent different evolutionary stages from a theorist’s point of view, for an observer the distinction could be important. The label ”WO” is a spectral classification, hence we keep using it here. As seen in Sect. 4, the detected features of I Zw 18 are consistent with WO stars.
@@
Nevertheless, the three phases do not differ much when it comes to the SED, yielding log Q values of 49.80 (early-pMS), 49.79 (mid-pMS) and 49.79 (late-pMS). Note how the surface temperatures of the evolutionary models are similar: the Teff of the early-pMS model is 138.9 kK, mid-pMS 139.2 kK, and late-pMS 138.0 kK (cf. Fig. 1 where the burning is color-coded). In terms of atmosphere models, Teff roughly reflects the wind onset region where 20; but even T2/3 (i.e. T at 2/3) is close for the three cases (with early-pMS 56.6 kK, mid-pMS 63.3 kK, late-pMS 70.3 kK), explaining why the predicted He II flux is not changing considerably during the core-helium burning phase. Thus, using late-pMS models in our population synthesis to stand for the entirety of the post main-sequence is well justified when it comes to the ionizing photon count.

### v3

@@ Abstract
Context. Chemically homogeneously evolving stars have been proposed to account for several exotic phenomena, including gravitational-wave emissions, gamma-ray bursts and certain types of supernovae.
@@
Results. Adding two new models computed with PoWR, we report that (i) these stars follow a unique sequence of classes: O WN WO (i.e. without ever being WC). From our population synthesis with standard assumptions, we predict that (ii) the source of the UV C IV 1550 Å and other emission bumps is a couple of dozen WO-type Wolf–Rayet stars (not WC as previously assumed) which are the result of chemically homogeneous evolution, while these, combined with the rest of the O-star population, account for the high He II ionizing flux and the spectral hardness. Contrasting our results against published optical and UV data from the literature and accounting for different aperture sizes and spatial regions probed by the observations, we find that (iii) our models are highly consistent with existing measurements.
Conclusions. Since our “massive Pop II stars” might just as well exist in early star-forming regions, our findings have implications for upcoming James Webb Space Telescope (JWST) surveys: the first galaxies in the high-redshift Universe may also experience the extra contribution of UV photons and the kinds of exotic explosions that chemically homogeneous stellar evolution predicts. Given that our results apply for binary populations too as long as the same fraction (10%) of the systems evolves chemically homogeneously, we conclude that the stellar progenitors of gravitational waves may very well exist today in I Zw 18.
@@
Galaxies: dwarf – Stars: massive – Stars: Wolf-Rayet – (Cosmology:) dark ages, reionization, first stars – Gravitational waves1 Introduction
@@
Dwarf starburst galaxies form stars at a high rate, and thus harbour large populations of metal-poor massive stars (e.g. Zhao:2013). Despite the recent burgeoning of interest in these stars’ lives (Weisz:2014; Izotov:2016; Schootemeijer:2018; Evans:2019; Garcia:2019; Groh:2019; Senchyna:2019; Franeck:2022; Szecsi:2022ok; Gull:2022; Lahen:2023) and deaths as exotic explosions (Perley:2016; AguileraDena:2018; Stevenson:2019; Agrawal:2022), and ‘afterlife’ as – potentially merging – compact objects (Eldridge:2016; Marchant:2016; deMink:2016; Mandel:2016; Marchant:2017; VignaGomez:2018; Eldridge:2019; Romagnolo:2022), theoretical models are far from being properly tested observationally (Mintz:2025).
@@
One convenient testbed for constraining our models is local dwarf star-forming galaxies. These objects typically display low metallicities and high star formation rates (Searle:1972; Hunter:1995; Vilchez:1998; Izotov:2002; Izotov:2004; Thuan:2005; Vaduvescu:2007; Shirazi:2012; Annibali:2013; Kehrig:2013; Kehrig:2018; Kehrig:2021; ArroyoPolonio:2024; Hirschauer:2024), meaning that they are local equivalents for high-redshift star-forming regions. Studying them can shed light on the nature of metal-poor massive-star populations that presumably contribute to the structure of the galaxy with their radiative, chemical, and mechanical feedback.
@@
In the blue compact dwarf galaxy I Zw 18 in particular, a high amount of photoionization has been derived from nebular line observations: Kehrig:2015 reported a He II ionizing flux of Q 1.331050 photons s-1 (see their Sect. 3.2). I Zw 18 has two large clusters: a north-western cluster and a south-western cluster. Of these, only the north-western cluster was spatially associated with this nebular emission; the total stellar mass of this region is about 300 000 M⊙ (Izotov:1998; Lebouteiller:2013). The high He II flux was puzzling because it could not be accounted for by the ionization from conventional ionizing sources including standard hot massive stars (Kehrig:2015).
@@
The hottest and most luminous massive stars are Wolf–Rayet (WR) stars: they are even stronger ionizing sources than O stars (Hamann:2015). Due to their optically thick stellar winds, WR stars also emit broad emission lines. One such line is the UV line C IV 1550 Å, which has indeed been identified in the spectrum of I Zw 18 (Lebouteiller:2013): in the north-western cluster, a line luminosity between L 2.25.51037 erg s-1 has been measured by Brown:2002. The problem is that according to classical stellar theory, the number of WR stars derived from the carbon-line luminosity cannot account for the high He II emission observed. As Kehrig:2015 report, to explain L (which they take to be to be 4.671037 erg s-1; see their Sect. 4.1), about nine classical Wolf–Rayet stars of type WC (with standard WC-properties given by Crowther:2006, see also LopezSanchez:2010a; LopezSanchez:2010b and Sander:2022IAU) are required in the region: and the ionizing flux of only nine of them is 50 times lower than Q.
@@
Therefore, other candidates were suggested as the source of ionization. One option is that a population of peculiar (nearly) metal-free massive stars (Pop III-star siblings/Pop III-like stars Kehrig:2015; Kehrig:2015b; Heap:2015) formed recently from some leftover primordial gas in this galaxy. Population III stars are expected to be hotter than metal-poor stars without developing carbon emission lines (Yoon:2012). This hypothesis relies on I Zw 18 keeping, or somehow obtaining, pockets of primordial gas that have just recently started to form massive stars.
@@
Another scenario was presented by Pequignot:2008, Lebouteiller:2017, and Schaerer:2019, in which a population of X-ray binaries produce the ionizing radiation. However, Kehrig:2021 found that neither a high-mass X-ray binary nor the soft X-ray photons observed in the galaxy can account for the bulk of the nebular HeII emission, and Senchyna:2020 support this conclusion. In turn, Oskinova:2022 suggest cluster winds and superbubbles as an alternative explanation.
@@
Here we study yet another possibility. We simulated stellar evolutionary models of single massive stars with the exact composition of I Zw 18 in Szecsi:2015, using state-of-the-art physics. We found that these “massive Pop-II stars” mostly evolve the classical way (similar to how metal-rich massive stars in our Galaxy do) as long as their rotational rate is moderate. But those models that rotate fast follow another path: they evolve chemically homogeneously, and become hot and luminous stars – similar to Wolf-Rayet stars but without the high wind-mass-loss and thus without the broad emission features during most of their lives. To analyse these objects further, we computed synthetic spectra for them in Kubatova:2019. We dubbed them transparent wind UV-intense stars (in short TWUIN stars) to distinguish them from traditional Wolf–Rayet stars with optically thick winds.
@@
Below we investigate whether a metal-poor massive star population based on our evolutionary models from Szecsi:2015 and our spectral models from Kubatova:2019 can simultaneously account for both the He II ionizing flux and the strength of emission lines observed in I Zw 18. Section 2 explains our modelling approach: from evolutionary models (Sect. 2.1) to synthetic spectra (Sect. 2.2, including the addition of new PoWR spectra and the newly established classification sequence for chemically homogeneously evolving stars) to population synthesis (Sect. 2.3) and emission line synthesis (Sect. 2.4). Section 3 compares our theoretical predictions to observations in terms of ionizing radiation and C IV 1550 Å emission. Then in Section 4 we dive into the existing literature to figure out if other emission lines observed so far in I Zw 18 (both optical and UV) support or contradict our theory. Section 5 discusses our explanation’s place amongst alternative ones as well as the necessary caveats. In Section 6 we conclude that, as far as the currently available observational data allows, I Zw 18 may – similarly to high-redshift galaxies (Liu:2025) – very well harbour chemically homogeneously evolving stars.
@@ 2.1 The models: Predicting O and WO stars
Metal-poor massive single stars have been modelled using the ‘Bonn’ stellar evolution code in Szecsi:2015 with the measured composition of I Zw 18 (totalling in a metallicity of Z=0.0002111Metallicities of dwarf galaxies are usually inferred from their nebular O III lines, which can be problematic since the ratio of iron to oxygen (and all other elements) does not necessarily follow solar patterns – on this, see Sect. 2.2.1 of Chruslinska:2019. Different works report different metallicities for I Zw 18, such as 1/40 Z⊙ in Kehrig:2016 and 1/30 Z⊙ in Izotov:1998; this may depend on, for example, the calibration used to translate O III equivalent width into abundances. Szecsi:2015 modelled I Zw 18 relying on element abundances reported in Lebouteiller:2013 corresponding to 1/50 Z⊙. , or [Fe/H] 1.7) and standard input physics. Mass-loss rate prescription from Hamann:1995 scaled down by a factor of 10 and Z-scaling from Vink:2001 was used for the hot helium-star phase, which yields consistent predictions with Nugis:2000. These evolutionary sequences predicted that fast-rotating stars evolve directly towards hot surface temperatures (so-called chemically homogeneous evolution, CHE) and thus serve as potential sources of photoionization in metal-poor dwarf galaxies. Applying simple estimations for the wind optical depth (following Langer:1989a) and black body radiation, Szecsi:2015 showed that during the long-lived main-sequence (core-hydrogen-burning) phase, these hot stars are expected to have transparent winds and therefore should not develop emission lines in their spectra (hence the name TWUIN stars). Szecsi:2015 estimated that a population of them with a regular mass function can explain the observationally derived He II ionizing flux, Q, in I Zw 18. However, these results were rather crude; a more reliable conclusion could be drawn from synthetic spectra created especially for these stars.
@@
In Kubatova:2019, we computed such spectra, again with the composition of I Zw 18, using the Potsdam Wolf-Rayet (PoWR) stellar atmosphere code (Grafener:2002; Hamann:2003; Hamann:2004; Surlan:2013; Sander:2015; Sander:2020). We found that chemically homogeneously evolving stars in the main-sequence phase indeed do not develop emission lines, just as suggested in Szecsi:2015. According to standard stellar classification presented in Figure 1, these TWUIN stars, if observed, would be classified as early O type stars (see also Table 4 & Appendix A of Kubatova:2019). As for the post-main-sequence, the evolutionary behaviour during this phase was not included into Szecsi:2015 but was later published by Szecsi:2022ok (their grid called ‘IZw18-CHE’) and also studied in Kubatova:2019. We found that during core-helium burning, chemically homogeneously evolving stars do actually develop strong emission lines: they were classified as WR stars of class WO. Fig. 1 summarizes these findings in a HR diagram, and Sect. 2.2 provides a deeper look into the post-main-sequence phase.
@@
The spectra published in Kubatova:2019 were created for three masses (Mini 20, 59 and 131 M⊙, all with a high initial rotational rate of 300-600 km-1), five evolutionary stages (four on the main-sequence and one post-main-sequence, see also Sect. 2.2), two mass-loss rates (a nominal one taken from the evolutionary models, and a reduced one; actual values listed in Table 2 of Kubatova:2019) and two types of assumptions about wind clumping (a smooth wind and a clumped wind with clumping factor D 10). The mass-loss rate values are supported by those that Hainich:2015 empirically predicted for the same stars. Additionally, more realistic mass-loss rates of these stars were calculated by Abdellaoui:2023 using the VH-1 code, which allows for a multi-dimensional treatment of the wind. The obtained results (listed in their Table 6.3) are somewhere between our nominal and reduced values.
@@
Fig. 2 shows the spectral energy distribution (SED) of some of our models (those with nominal mass-loss rates and smooth wind), as well as the corresponding black body spectrum for context (without applying wind optical depth correction from Langer:1989a). As Fig. 2 attests, the SED of these stars do not follow the distribution of black body radiation, even though the total amount of ionizing photons in the three main ionizing continua (Lyman, He I and He II continua) are on the same order of magnitude as in the (uncorrected) black body.
@@
Having these models allows us to build a synthetic population that includes spectral synthesis of various emission lines. Before we do that (Sect. 2.3), we present two new model spectra to properly account for the post-main-sequence, a phase crucially relevant for our goals.
@@ 2.2 New PoWR spectra: WN and WO stars (no WC)
In our Kubatova:2019, the post-main-sequence phase was accounted for by models with central helium mass fraction of YC 0.1; that is, nearly the end of core-helium burning. The spectra were classified as type WO according to the standard spectral classification scheme (for details on this, see Table 4 & Appendix A of Kubatova:2019, as well as our Fig. 1), and showed very prominent emission lines in both the optical and the UV regimes. However, according to Figure 3, the evolutionary model predicts significant change in the surface abundances while core-helium burning is in progress, with the YC 0.1 stage being quite different from earlier times. Therefore, before using these late-stage models in our population synthesis (Sect. 2.3), we should justify the extent to which they are representative of the entirety of the post-main-sequence, and discuss their evolutionary behaviour and spectral classification.
@@
To this end, we computed two more spectral models using PoWR for YC 0.8 and YC 0.5 of a very massive star with initial mass Mini 131 M⊙ (actual mass around 93106 M⊙, see the caption). The computational set-up was the same as in Kubatova:2019. We apply nominal mass-loss rates taken from the evolutionary model (Szecsi:2015; Szecsi:2016; Szecsi:2022ok), and smooth winds (clumping factor D 1); we discuss in Sect. 5.2 that these assumptions are actually quite robust. Figure 4 compares the optical band for these spectra, while Figure 3 summarizes what we learned in terms of stellar classification.
@@
Applying the rudimentary scheme of classification defined by Georgy:2012 for evolutionary models, Szecsi:2016 suggested that these stars are of the classes WN, WC, and close-to-WO, respectively (Fig. 3). However, the PoWR spectra tell a different story. The standard classification scheme (working with spectral line intensity222This classification scheme follows the footsteps of the Harvard-system. The Harvard system actually included WR-stars, calling them Oa (now WC), Ob (now WNE), Oc (now WNL) – see Cannon:1916 – but efforts like Payne:1930, Edlen:1932, and Beals:1933 helped shape the version adopted by the IAU in 1938, which was modernized in the 1960s by Hiltner, Schild and Lindsey Smith, with WO stars first introduced by Barlow:1982. Today, we rely on emission line intensities to assign classes to WR-stars, as summarized in Appendix A of Kubatova:2019, although switching between subtypes is not as clear as for the absorption-line-based classes OBAFGKM and the usage of specific primary and secondary criteria is also less sharp. as summarized by Appendix A of Kubatova:2019) yields the following classes: WN2 for YC 0.8, WO 2 or WO 1 (consistent with both) for YC 0.5, and WO 1 for YC 0.1. Based on what we learned in Kubatova:2019 (especially their Figs. 4 and B6), we expect that reducing the mass loss rate and/or increasing the wind clumping may shift these spectra towards different subtypes – that is, towards different line intensities – but not towards a different class. In other words, varying mass loss and clumping will not make any of these models a WC star.
@@
We conclude therefore that chemically homogeneously evolving stars at I Zw 18 metallicity would be seen, after spending their main-sequence as O stars, as early-type WN stars at the beginning, and WO stars in the majority of their post-main-sequence lifetimes. This is surprising since typical evolution calculations for single stars predict the WO stage to be a rather late, short-term phase, preceded by a lengthier WC phase (see e.g. Groh:2014), the typical sequence of classes reading:
O RSG/LBV WN WC WO.
@@
Chemically homogeneous evolution does not follow that picture, as these stars complete their main-sequence as very hot O stars (Sect. 2.1 and Fig. 1) and then proceed to be WN and WO after that, a completely new sequence of classes:
O WN WO.
@@
According to Fig.3, if a WC-phase arises at all between WN and WO, it will not last longer than about 1-2% of the total stellar lifetime.
@@
Finding our stars to be WO instead of WC is in accordance with Tramper:2013 who concluded that WO stars may be the high-temperature and high-luminosity extensions of WC stars. The WO spectral type does not necessarily mean more oxygen than carbon: in our mid-pMS model, the carbon abundance is actually higher than the oxygen one (as is seen in Fig. 3). The same was found for the LMC WO stars analysed in Aadland:2022. So it is not about abundance but about temperature: carbon is ionized into C V and thus the otherwise prominent 1550 and 5808 features are weak or absent.
@@
While WC and WO may not necessarily represent different evolutionary stages from a theorist’s point of view, for an observer the distinction could be important. The label “WO” is a spectral classification, hence we keep using it here. As is seen in Sect. 4, the detected features of I Zw 18 are consistent with WO stars.
@@
Nevertheless, the three phases do not differ much when it comes to the SED, yielding log Q values of 49.80 (early-pMS), 49.79 (mid-pMS) and 49.79 (late-pMS). Note how the surface temperatures of the evolutionary models are similar: the Teff of the early-pMS model is 138.9 kK, mid-pMS 139.2 kK, and late-pMS 138.0 kK (see Fig. 1 where the burning is colour-coded). In terms of atmosphere models, Teff roughly reflects the wind onset region where 20; but even T2/3 (i.e. T at 2/3) is close for the three cases (with early-pMS 56.6 kK, mid-pMS 63.3 kK, late-pMS 70.3 kK), explaining why the predicted He II flux is not changing considerably during the core-helium burning phase. Thus, using late-pMS models in our population synthesis to stand for the entirety of the post-main-sequence is well justified when it comes to the ionizing photon count.